# MIMIR: Masked Image Modeling for Mutual Information-based Adversarial Robustness

### v3

@@ Abstract
Vision Transformers (ViTs) achieve excellent performance in various tasks, but they are also vulnerable to adversarial attacks. Building robust ViTs is highly dependent on dedicated Adversarial Training (AT) strategies. However, current ViTs’ adversarial training only employs well-established training approaches from convolutional neural network (CNN) training, where pre-training provides the basis for AT fine-tuning with the additional help of tailored data augmentations. In this paper, we take a closer look at the adversarial robustness of ViTs by providing a novel theoretical Mutual Information (MI) analysis in its autoencoder-based self-supervised pre-training. Specifically, we show that MI between the adversarial example and its latent representation in ViT-based autoencoders should be constrained by utilizing the MI bounds. Based on this finding, we propose a masked autoencoder-based pre-training method, MIMIR, that employs an MI penalty to facilitate the adversarial training of ViTs. Extensive experiments show that MIMIR outperforms state-of-the-art adversarially trained ViTs on benchmark datasets with higher natural and robust accuracy, indicating that ViTs can substantially benefit from exploiting MI. In addition, we consider two adaptive attacks by assuming that the adversary is aware of the MIMIR design, which further verifies the provided robustness. Our code and trained models are publicly available at the anonymous link. [^1] https://anonymous.4open.science/r/MIMIR-5444
1 Introduction
Adversarial training (Madry et al. 2018) is one of the standard techniques to build robust models. Regarding adversarial training for ViTs, early works focus on investigating the unique attention mechanism of ViTs. There, robustness can be improved by randomly dropping gradient according to attention (Mo et al. 2022) or improving training efficiency by dropping low-attention image embeddings (Wu et al. 2022). The majority of recent works focus on improving the ViTs’ adversarial training based on experience with CNNs, given AT’s success in building robust CNNs. Looking at the adversarial training recipe differences between ViTs and CNNs, strong data augmentation, e.g., Randaugment (Cubuk et al. 2020), CutMix (Yun et al. 2019), and MixUp (Zhang et al. 2018), show positive influence in CNN but are ineffective for ViTs AT and sometimes even reduce the AT performance of ViTs (Bai et al. 2021; Singh, Croce, and Hein 2023). To mitigate this, the distortion magnitudes of Randaugment and the probability of applying MixUp or CutMix need to be progressively increased (Bai et al. 2021). Similar problems of strong data augmentations are also observed by (Debenedetti, Sehwag, and Mittal 2023), where basic data augmentations, i.e., random-resize-and-crop, horizontal-flipping, and color-jitter with high weight decay are shown to be promising to improve ViT adversarial training.
Another common practice in ViT adversarial training is to leverage pre-trained models, which has also been extensively explored (Mo et al. 2022; Singh, Croce, and Hein 2023). Fine-tuning on natural pre-trained models could benefit the adversarial robustness of ViTs. For instance, recently, AdvXL (Wang et al. 2024) introduced an efficient and effective adversarial training strategy for giant models on web-scale datasets. However, as the current ViT adversarial training strategies are mainly inherited from CNN AT, the key factor that drives the success of pre-training has not been completely revealed. For instance, pre-training with ImageNet-21K and SimMIM (Xie et al. 2022) performs similarly to the model trained from scratch, and even the CLIP (Radford et al. 2021) pre-trained model shows a performance drop (Liu et al. 2024).
In this work, we aim to provide an in-depth analysis of ViT’s robustness using MI, focusing on self-supervised pre-training. Information Bottleneck (IB) has been previously used in supervised learning to analyze adversarial robustness, where a joint objective of simultaneously minimizing the MI between inputs and latent features while maximizing the MI between labels and latent features is introduced to mitigate the impact of the adversarial noise in the inputs. Regarding the task of ViT adversarial training, we provide a novel theoretical analysis for the self-supervised autoencoder, showing that the self-supervised ViT benefits from decreasing the MI between inputs and latent features. Based on this finding, we propose a theoretically grounded adversarial pre-training method, Masked Image Modeling for Mutual Information-based Adversarial Robustness (MIMIR). Figure 1 provides an illustrative diagram of MIMIR.
MIMIR substantially improves the adversarial robustness of ViT models as demonstrated by comprehensive experiments on ImageNet-1K (Deng et al. 2009), Tiny-ImageNet (Le and Yang 2015), and CIFAR-10 (Krizhevsky, Hinton et al. 2009). We also provide an experimental comparison with the state-of-the-art, showing the effectiveness of MIMIR in improving both natural and adversarial prediction accuracy (see Section 4.2).
Our main contributions are:
@@ 1 Introduction
  • By revising the current ViT AT strategies, we point out that ViT adversarial pre-training methods compromise natural accuracy and lack systematic analysis. To this end, we provide a theoretical analysis using adversarial examples and the Mutual Information penalty. The MI bounds motivate us to decrease MI between adversarial examples and learned latent representation.
@@
  • Based on the findings of our analysis, we propose an efficient self-supervised defense MIMIR against adversarial attacks on ViTs. We systematically evaluate MIMIR using three datasets and multiple model scales under various adversarial attacks, demonstrating its superiority.
  • We show MIMIR is resistant to two adaptive attacks, where the adversary is aware of MIMIR design. We build a PGD-fea attack that increases the distance between features extracted from natural and adversarial examples. Moreover, we build a PGD-MI attack that includes the MI penalty as the learning objective to find adversarial examples.
@@
Figure 1: Diagram illustrating the working mechanism of MIMIR. In the pre-training, adversarial perturbations $\delta$ are generated and added to natural images $x$ to create adversarial images $x+\delta$, as shown on the left. Patches of generated adversarial images are fed as training data to the ViT-based autoencoder, where the output of decoder $x_{re}$ and the natural input image $x$ are used to calculate the loss. In particular, we propose using MI ($I(x+\delta,z)$) as an additional penalty to compose the pre-training loss of MIMIR, as shown in the bottom right. After pre-training, a trained encoder is combined with a randomly initialized classification layer as the final model that is further fine-tuned for classification tasks.
2 Background and Related Work
[…]
@@ 2 Background and Related Work
The generated adversarial examples are as similar as possible to the original natural examples by limiting $\delta$ to a relatively small domain:
[…]
where $S$ is the $l_{\infty}$-ball of radius $r$ at a position $x_{i}$ in $\mathbb{R}^{d_{X}}$. The distance between $x_{i}$ and $x_{i}+\delta$ can be evaluated by norms such as $l_{0},l_{2},$ and $l_{\infty}$.
When using the above adversarial examples to conduct adversarial training, the learning objective is:
@@
Information Bottleneck (IB). The IB concepts were first proposed in (Tishby, Pereira, and Bialek 2000) and further developed for deep learning in (Tishby and Zaslavsky 2015; Shwartz-Ziv and Tishby 2017). It has been widely used to improve adversarial robustness (Xu, Perin, and Picek 2023; Alemi et al. 2017; Yu, Yu, and Príncipe 2021; Wang et al. 2021). IB describes the generalization of a deep neural network in two phases: 1) empirical error minimization (ERM) and 2) representation compression. For a network with input $x$ and label $y$, there is an intermediate representation $t_{l}$ for each layer $l$, i.e., the output of $l$-th layer. The IB principle aims to keep more relevant information in $t_{l}$ about target $y$ while decreasing the irrelevant information about input $x$. The information between intermediate representation $t_{l}$ and input $x$ or label $y$ is quantified by MI, denoted by $I$.
Adversarial Robustness of ViTs. PGD (Madry et al. 2018) adversarial training is considered one of the most effective defenses that can withstand adaptively designed attacks (Athalye, Carlini, and Wagner 2018). Due to the difference between CNNs and ViTs, there have been substantial recent efforts to explore new adversarial training approaches for ViTs. EasyRobust (Mao et al. 2022) evaluates the robustness of ViTs using PGD adversarial training. (Mo et al. 2022; Bai et al. 2021; Debenedetti, Sehwag, and Mittal 2023) demonstrate the classical training recipe for CNN is not suitable for ViTs and present new adversarial training strategies for ViTs by evaluating different hyperparameters and data augmentation policies. (Mo et al. 2022) also finds that pre-training with natural data provides better robustness after adversarial fine-tuning. (Singh, Croce, and Hein 2023) explores the effect of strong pre-training and replacing the patch embedding in ViTs with convolutional blocks. AdvXL (Wang et al. 2024) shows an outstanding performance of giant models and web-scale data with adversarial pre-training and fine-tuning. However, (Liu et al. 2024) provides a different view that pre-training with ImageNet-21K and SimMIM (Xie et al. 2022) performs similarly to the model trained from scratch, and the CLIP pre-trained model has a performance drop. Unlike existing works, we provide a formal and in-depth analysis leveraging MI, based on which we introduce a new adversarial pre-training scheme.
3 MIMIR
3.1 Threat Model
Adversary’s Goal. The attacker aims to fool the trained model by decreasing the overall classification accuracy (non-targeted) or forcing the model to recognize an input as a specific target (targeted). The image perturbations are limited by the $l_{\infty}$ norm for invisibility.
@@ 3.1 Threat Model
Adversary Knowledge. The attacker has white-box access to the model, including training data, architectures, hyperparameters, and model weights. The attacker can implement iterative attacks and unlimited queries to update adversarial examples multiple times in white and black-box settings. Adversarial examples can be created according to model architectures, parameters, the gradients of loss function, and datasets. The adversary can adaptively design new attacks for a specific model according to the design details of the defense method.
Defender’s Goal. The defender’s main goal is to train a robust model against potential adversarial attacks and prevent the performance of natural data from decreasing significantly. In addition, the defense method should be robust against adaptive attacks specifically designed for it.
3.2 Design Intuition
MIM has been proved useful as a pre-training method for ViTs on various tasks (Vincent et al. 2010; He et al. 2022; Wei et al. 2022; Bao et al. 2022). To train a powerful model using MIM, one can mask out a part of the foreground of inputs and reconstruct it using the model. Masking out the foreground instead of the background reduces discriminative information in visible information to the model. Reconstructing foreground parts is harder than background and helps the model learn more discriminative information (Wang et al. 2023). Inspired by this phenomenon, we aim to build a more difficult task by adding adversarial perturbations to natural inputs. The adversarial perturbations increase the distance between the input and reconstruction target. If we can reconstruct natural inputs from adversarial examples, it means the features learned by the model are robust against adversarial attacks. In other words, we want the encoder to learn a latent representation that does not carry information concerning adversarial perturbations while enabling the decoder to reconstruct the natural image. Note that simply masking out all foreground does not make sense, as it would render the reconstruction of meaningful content infeasible. From the perspective of IB, there is a bottleneck between the encoder and decoder. Our adversarial pre-training task contains two information sources: the natural data $x$ and the adversarial perturbations $\delta$. As the information flows through the bottleneck, adversarial information from $\delta$ is eliminated. The natural information from $x$ is maintained because of the constraint when reconstructing the target $x$.

### v5

@@ Abstract
Vision Transformers (ViTs) have emerged as a fundamental architecture and serve as the backbone of modern vision-language models. Despite their impressive performance, ViTs exhibit notable vulnerability to evasion attacks, necessitating the development of specialized Adversarial Training (AT) strategies tailored to their unique architecture. While a direct solution might involve applying existing AT methods to ViTs, our analysis reveals significant incompatibilities, particularly with state-of-the-art (SOTA) approaches such as Generalist [1] (CVPR 2023) and DBAT [2] (USENIX Security 2024). This paper presents a systematic investigation of adversarial robustness in ViTs and provides a novel theoretical Mutual Information (MI) analysis in its autoencoder-based self-supervised pre-training. Specifically, we show that MI between the adversarial example and its latent representation in ViT-based autoencoders should be constrained via derived MI bounds. Building on this insight, we propose a self-supervised AT method, MIMIR, that employs an MI penalty to facilitate adversarial pre-training by masked image modeling with autoencoders. Extensive experiments on CIFAR-10, Tiny-ImageNet, and ImageNet-1K show that MIMIR can consistently provide improved natural and robust accuracy, where MIMIR outperforms SOTA AT results on ImageNet-1K. Notably, MIMIR demonstrates superior robustness against unforeseen attacks and common corruption data and can also withstand adaptive attacks where the adversary possesses full knowledge of the defense mechanism. Our code and trained models are publicly available at: https://github.com/xiaoyunxxy/MIMIR.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23 - 27 February 2026 , San Diego, CA, USA ISBN 979-8-9919276-8-0 https://dx.doi.org/10.14722/ndss.2026.241813 www.ndss-symposium.org
I Introduction
ViTs [3] and their variants [4, 5] have achieved substantial progress and serve as foundational components in modern vision-language models. Prominent multimodal frameworks, including CLIP [6], BLIP [7], and Mini-GPT4 [8], typically employ ViTs as their image encoders. However, similar to convolutional neural networks (CNNs), attention-based models provide limited robustness against evasion attacks [9, 10, 11, 12, 13]. Evasion attacks [14, 15] (also known as adversarial attacks), where well-trained deep models are fooled by introducing human-imperceptible perturbations to inputs, remain a persistent challenge in deep learning security. In 2024, the National Institute of Standards and Technology (NIST) explicitly listed adversarial attacks as a significant threat to AI systems, and pointed out the importance of conducting robustness testing and mitigation, such as adversarial training and formal verification, when deploying AI tools [16]. Nevertheless, improving adversarial robustness remains a difficult task, where even SOTA methods, such as [17, 18, 19], achieve only marginal robustness gains, commonly below 2% compared with those before them.
So far, Adversarial Training (AT) is widely recognized as the most practically effective defense [12, 11, 20] against evasion attacks. AT operates by augmenting the training dataset with adversarially perturbed samples [21], yet introduces two key limitations: (1) substantial computational overhead due to the generation of adversarial examples during training [21], and (2) a potential degradation in natural accuracy [22]. Numerous methods have been proposed to mitigate these challenges. Techniques such as FreeAT [23] optimize efficiency by reusing gradient information during adversarial example generation, while FastAT [24] employs an improved Fast Gradient Sign Method (FGSM) to accelerate training. TRADES [25], SCORE [26], Generalist [1], and DBAT [2] explore how to achieve the best trade-off between natural and robust accuracy. Additionally, pre-training strategies have also been leveraged to enhance the performance of AT [27, 17].
Applying existing AT methods to ViTs presents unique challenges due to the fundamental differences between attention-based architectures and CNNs. Unlike CNNs, ViTs lack inductive biases [3], including locality, two-dimensional neighborhood structure, and translation equivariance. These biases are inherent to CNNs as prior knowledge, enabling efficient learning with limited data, whereas ViTs typically require larger training datasets to achieve comparable generalization performance [3]. Consequently, AT for ViTs entails significantly higher computational costs. Initial research on AT for ViTs explored their unique attention mechanism. For instance, robustness can be improved by randomly dropping gradients according to attention [11] or improving training efficiency by dropping low-attention image embeddings [12]. The majority of recent works have focused on adapting CNN-based AT methodologies to ViTs, given AT’s success in building robust CNNs. Unfortunately, standard CNN AT techniques are not fully transferable to ViTs. Empirical studies [9, 20] reveal that strong data augmentations (such as Randaugment [28], CutMix [29], and MixUp [30], which improve robustness in CNNs) often degrade AT performance for ViTs. To mitigate this, recent work [9] suggests progressively increasing augmentation intensity (e.g., distortion magnitudes in RandAugment or the sampling probability of MixUp/CutMix) during training. Furthermore, SOTA AT strategies, such as Generalist [1] and DBAT [2], are less effective when applied to ViTs (see Table I), and there is a lack of evaluation on large datasets, such as ImageNet-1K, which further limits their generalizability.
Pre-training has emerged as a complementary approach to ViT AT, with studies showing that adversarial fine-tuning of naturally pre-trained models can enhance robustness [11, 17]. AdvXL [31] notably advanced this paradigm by developing efficient AT for web-scale datasets. However, the mechanisms underlying pre-training’s effectiveness are not fully understood, and results are inconsistent across implementations. For instance, models pre-trained on ImageNet-21K using SimMIM [32] demonstrate comparable performance to scratch-trained counterparts, while CLIP [6] pre-training has been shown to degrade performance in some configurations [33].
While previous methods of ViT AT, such as [17, 20, 9], focus on searching for better combinations of training hyperparameters, they suffer from performance drops across different architectures and datasets. In contrast, we aim for a generalizable method via pre-training. Specifically, this work presents a systematic investigation of self-supervised pre-training for ViT robustness through the lens of Mutual Information (MI) and Information Bottleneck (IB) theory. IB introduces a joint objective of simultaneously minimizing the MI between inputs and latent features while maximizing the MI between labels and latent features to mitigate the impact of the adversarial noise in the inputs. Regarding the ViT AT, we develop a novel theoretical justification for self-supervised autoencoders, demonstrating that reducing MI between inputs and latent features enhances ViT robustness. Based on this finding, we propose a theoretically grounded adversarial pre-training method, Masked Image Modeling for Mutual Information-based Adversarial Robustness (MIMIR).  [^1] Mimir is a figure in Norse mythology, renowned for his knowledge and wisdom. Specifically, we convert Masked Image Modeling (MIM) into an effective and efficient adversarial pre-training method. The basic idea is to predict the masked content of inputs, which is a self-supervised learning task. The effectiveness of MIMIR is analyzed and guaranteed by our theoretical justification. The efficiency comes from discarding the masked content (75% of image patches are discarded in our experiments), which greatly reduces the computing requirements. Figure 1 provides an illustrative diagram of MIMIR.
We validate MIMIR’s effectiveness through extensive experiments on CIFAR-10 [34], Tiny-ImageNet [35], and ImageNet-1K [36], showing consistent improvements in both natural and adversarial accuracy. In addition, we test the generalizability of MIMIR by combining MIMIR with three MIM methods for three representative architectures, including MAE [37] for ViTs, Group Window Attention [38] for hierarchical transformer (Swin [39]), and SparK [40] for CNN (ConvNext [41]), where MIMIR outperforms SOTA AT methods on ImageNet-1K. Our main contributions are:
@@ I Introduction
  • By revisiting the current ViT AT strategies, we point out that ViT AT methods compromise natural accuracy and lack systematic study. To this end, we provide a theoretical analysis using adversarial examples and the Information Bottleneck. We theoretically show that the Mutual Information between adversarial examples and the learned latent representation should be decreased for better robustness.
@@
  • Based on our analysis, we propose a self-supervised defense, MIMIR, against adversarial attacks on ViTs. We evaluate MIMIR using multiple architectures on three datasets under various adversarial attacks, demonstrating its effectiveness. MIMIR achieves SOTA adversarial robustness on ImageNet-1K following the standardized evaluation by RobustBench. [^2] https://robustbench.github.io/
  • We show that MIMIR is robust against unforeseen attacks and common corrupted data (ImageNet-C [42]) and can resist adaptive attacks where the adversary is aware of MIMIR’s design.
@@
Figure 1: Diagram illustrating the working mechanism of MIMIR. In the pre-training, adversarial images $x+\delta$ are generated and separated into image patches as the inputs of the encoder. The output of decoder $x_{re}$ and the natural input image $x$ are used to calculate the loss. After pre-training, a trained encoder is combined with a randomly initialized classification layer as the final complete model. Then the complete model is further fine-tuned for classification tasks.
II Background
II-A Evasion Attack
Evasion attacks [43, 14, 15], also known as adversarial attacks, refer to applying imperceptible perturbations to the original input of the machine learning model (in this work, neural network), which generates adversarial examples to fool the victim model. Given an $L$-layer neural network $F_{\theta}$ for classification in $d_{Y}$-dimensional space and a training dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ in $d_{X}$-dimensional space, the two primary goals of adversarial attacks are:
  1. 1.
    The generated perturbation $\delta$ can successfully mislead the network by maximizing (e.g., PGD attack [21]):
@@ II-A Evasion Attack
    where $x_{i}\in\mathbb{R}^{d_{X}}$ and $y_{i}\in\{0,1\}^{d_{Y}}$, $\theta$ are the parameters of the current network and $\mathcal{L}_{CE}$ is the standard CE (Cross-Entropy) loss. The perturbation aims to decrease confidence in the ground truth labels while increasing confidence in wrong labels. Thus, the loss between the misleading outputs and the ground truth labels increases.
  2. 2.
    The generated adversarial examples are as similar as possible to the original clean examples by limiting $\delta$ to a relatively small domain:
    $S=B(x_{i},r)=\{\delta\in\mathbb{R}^{d_{X}}:\left\|\delta\right\|_{\infty}\leq r\},$ (2)
    where $S$ is the $l_{\infty}$-ball of radius $r$ at a position $x_{i}$ in $\mathbb{R}^{d_{X}}$. The distance between $x_{i}$ and $x_{i}+\delta$ can be evaluated by norms such as $l_{2}$ and $l_{\infty}$.
When using the above attacks to generate adversarial examples for AT, the learning objective is:
@@
II-B Masked Image Modeling - MIM
MIM refers to a self-supervised pre-training framework that aims to reconstruct pre-defined targets, such as discrete tokens [44], raw RGB pixels [37, 45], or features [46]. The final goal is to use the pre-trained model as a starting point for downstream fine-tuning. The downstream tasks include, for instance, classification and object detection. More specifically, to build a high-performance ViT $f_{e}$ without a classification layer, we consider $f_{e}$ as an encoder to extract discriminative input features. Then, we design a lightweight decoder $f_{d}$, which uses the output of $f_{e}$ as its input. The goal of $f_{d}$ is to reconstruct the original inputs (let us consider MAE [37] as an example). The aim is to decrease the distance between the input $x$ and $x_{re}=f_{d}\circ f_{e}(x)$. After the encoder $f_{e}$ and decoder $f_{d}$ are trained, we use $f_{e}$ plus a manually initialized classification layer as the starting point of fine-tuning.
II-C Mutual Information - MI
MI measures the mutual dependence between two random variables, $X$ and $Y$. It quantifies the amount of information contained in one random variable about another random variable or the reduced uncertainty of a random variable when another random variable is known. It can be written as:
$I(X,Y)=\int_{\mathcal{Y}}\int_{\mathcal{X}}P_{(X,Y)}(x,y)\log\left(\frac{P_{(X,Y)}(x,y)}{P_{(X)}(x)P_{(Y)}(y)}\right),$ (4)
where $P_{(X,Y)}$ is the joint probability density function of $X$ and $Y$. $P_{(X)}$ and $P_{(Y)}$ are the marginal probability density function of $X$ and $Y$. MI can be equivalently expressed as:
$I(X,Y)$ $=H(X)-H(X|Y).$ (5)
Estimating the exact MI is not easy, as it is difficult to precisely estimate $P_{X,Y}$ or $P_{X}$ and $P_{Y}$ in high-dimensional space [47]. In practice, Deep Deterministic Information Bottleneck (DIB) [48] suggested using the matrix-based Renyi’s $\alpha$-order entropy $I_{\alpha}$ [49, 50] to estimate MI, which avoids density estimation and variational approximation. An alternative way is the Hilbert-Schmidt Independence Criterion (HSIC) [51], which is a kernel-based dependence measure defined in a reproducing kernel Hilbert space (RKHS) and is usually used as a surrogate of MI. Details about definitions and empirical estimators of $I_{\alpha}$ and HSIC are provided in Appendix -J. In this paper, we evaluate both $I_{\alpha}$ and HSIC as our MI measurements.
II-D Information Bottleneck - IB
@@
The IB concepts were first proposed in [52] and further developed for deep learning in [53, 54]. IB describes the generalization of a deep network in two phases: 1) empirical error minimization (ERM) and 2) representation compression [54]. For a network with input $x$ and label $y$, there is an intermediate representation $t_{l}$ for each layer $l$, i.e., the output of the $l$-th layer. The IB principle aims to keep more relevant information in $t_{l}$ about the target $y$ while decreasing the irrelevant information about the input $x$. The information between intermediate representation $t_{l}$ and input $x$ or label $y$ is quantified by MI, denoted by $I(\cdot)$. During neural network training, in the ERM phase, the model increases shared information between $t_{l}$ with respect to both $x$ and $y$. Afterward, in the compression phase, the model decreases information contained in $t_{l}$ about $x$ but preserves (or even increases) information about $y$. The reduction of $I(x,t_{l})$ can be interpreted as a way of reducing noise or compressing irrelevant or redundant features in $x$. At the end of the training, the model strikes a trade-off that maximizes $I(y,t_{l})$ and minimizes $I(x,t_{l})$. Formally, the IB minimizes the following Lagrangian:
$\mathcal{L}=I(x,t_{l})-\beta I(y,t_{l}),$ (6)
where $\beta$ is a Lagrange multiplier that controls the trade-off between predicting $y$ and compressing $x$.
III MIMIR
III-A Threat Model
Adversary’s goal. The attacker aims to fool the trained model with both non-targeted and targeted attacks. The goal is to decrease the overall classification accuracy (non-targeted) or compel the model to recognize any inputs as a specific target (targeted). Meanwhile, the adversarial perturbations applied to the input should be invisible so that they will not be easily detected. During the training phase, the model optimizes its parameters to minimize the loss between predicted outputs and true labels, thereby enhancing classification accuracy. In contrast, the adversary’s objective is to develop algorithms that generate perturbations capable of maximizing this loss. For a targeted attack, the attacker decreases the loss between the output and the specified target label. To maintain the imperceptibility of the perturbations, the magnitude of the adversarial modifications is constrained by distance measures (such as $l2$ and $l_{\infty}$), ensuring that the alterations to the input data remain within a visually indistinguishable range.
@@ III-A Threat Model
Adversary knowledge. The attacker has white-box access to the model, including training data, architectures, hyperparameters, and model weights. The attacker can implement iterative attacks and unlimited queries to update adversarial examples multiple times in white and black-box settings. Adversarial examples can be created according to model architectures, parameters, the gradients of the loss function, and datasets. In addition, we also consider adaptive adversaries who are also aware of potential defenses. The adversary can design new attacks for a specific model according to the design details of the defense method.
Defender’s goal. From the defender’s perspective, the main goal is to train a robust model against potential adversarial attacks. The defender considers the following objectives:
  • The defender aims to prevent the performance of natural data from decreasing significantly, but allows a slight drop of natural accuracy for a trade-off in exchange for robustness.
  • The defense method should provide a notable improvement compared to models without defenses when subjected to various adversarial attacks, especially to adaptive attacks that are aware of the details of the defense method.
  • The defense method should be efficient and scalable to large datasets such as ImageNet-1K [36].
III-B Design Intuition
MIM has been established as an effective pre-training paradigm for Vision Transformers (ViTs), demonstrating strong performance across diverse downstream tasks [55, 37, 46, 44]. The core methodology involves masking foreground regions of input images and tasking the model with their reconstruction. Masking foreground (as opposed to background) regions removes high-information content and results in a harder task (than reconstructing background content), forcing the model to develop stronger feature representations to reconstruct semantically meaningful patterns [56].
Building upon these principles, we introduce an adversarial extension to MIM by incorporating adversarial perturbations into the input space. Our formulation is grounded in three interconnected hypotheses:
  • Adversarial Robustness through Reconstruction: If a model can reconstruct natural images from adversarially perturbed inputs, its latent representations must inherently discard perturbation-specific information while preserving natural data semantics.
  • IB Perspective: The encoder-decoder architecture naturally imposes an information bottleneck. When processing adversarial examples $x+\delta$, the system must suppress perturbation-derived information $(\delta)$ while retaining natural data information $(x)$ to achieve accurate reconstruction (see Figure 1).
  • Optimal Masking Strategy: Complete foreground masking (to build a difficult task) is suboptimal, as it eliminates essential reconstruction signals. Instead, our method employs partial masking to maintain a tractable yet challenging learning objective.